OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (2, 5) and (–1, 8)?

OpenStudy (anonymous):

y-5=-1(x-2)

OpenStudy (cgreenwade2000):

I'm sorry but that is incorrect.

OpenStudy (anonymous):

???

OpenStudy (cgreenwade2000):

For point slope form you need to find the slope. Can you tell me how to find the slope using two points?

OpenStudy (anonymous):

use this formula slope = (y2-y1)/(x2-x1)

OpenStudy (anonymous):

jeyde205 did u got the ans

OpenStudy (anonymous):

yeahh thanks

OpenStudy (anonymous):

ok